# PointSAM: Pointly-Supervised Segment Anything Model for Remote Sensing Images

### v1

@@ Abstract
Segment Anything Model (SAM) is an advanced foundational model for image segmentation, widely applied to remote sensing images (RSIs). Due to the domain gap between RSIs and natural images, traditional methods typically use SAM as a source pre-trained model and fine-tune it with fully supervised masks. Unlike these methods, our work focuses on fine-tuning SAM using more convenient and challenging point annotations. Leveraging SAM’s zero-shot capabilities, we adopt a self-training framework that iteratively generates pseudo-labels for training. However, if the pseudo-labels contain noisy labels, there is a risk of error accumulation. To address this issue, we extract target prototypes from the target dataset and use the Hungarian algorithm to match them with prediction prototypes, preventing the model from learning in the wrong direction. Additionally, due to the complex backgrounds and dense distribution of objects in RSI, using point prompts may result in multiple objects being recognized as one. To solve this problem, we propose a negative prompt calibration method based on the non-overlapping nature of instance masks. In brief, we use the prompts of overlapping masks as corresponding negative signals, resulting in refined masks. Combining the above methods, we propose a novel Pointly-supervised Segment Anything Model named PointSAM. We conduct experiments on RSI datasets, including WHU, HRSID, and NWPU VHR-10, and the results show that our method significantly outperforms direct testing with SAM, SAM2, and other comparison methods. Furthermore, we introduce PointSAM as a point-to-box converter and achieve encouraging results, suggesting that this method can be extended to other point-supervised tasks. The code is available at https://github.com/Lans1ng/PointSAM.
@@ I Introduction
Foundation models are versatile, large-scale models designed for a wide range of tasks and applications. They have demonstrated exceptional performance in areas such as natural language processing (e.g., BERT[1] and GPT-3[2]) and multimodal tasks (e.g., CLIP[3] and ALIGN [4]). Recently, the Segment Anything Model (SAM) [5, 6] was introduced as a foundation model specifically for image segmentation. Trained on a billion-scale dataset of masks and prompts, SAM can be applied to various downstream tasks requiring promptable segmentation, including healthcare [7, 8], autonomous driving [9], and remote sensing [10, 11, 12].
Despite SAM’s strong zero-shot capabilities, challenges persist in handling out-of-distribution (OOD) data and domain shifts in remote sensing images (RSIs). Many categories in RSIs are not represented in SAM’s training data, and RSIs, typically captured from aerial or satellite perspectives, differ significantly from natural images. Consequently, recent studies [10, 13, 14, 15] have focused on fine-tuning SAM for specific tasks. For example, RS-Prompter [10] uses queries or anchors as prompts to guide SAM’s mask decoder for instance segmentation. Similarly, SAM-CD [13] employs FastSAM’s encoder and introduces adapters for fine-tuning in change detection tasks.
@@
While these methods have shown promising results, they also require full mask annotations, which are difficult and time-consuming to obtain. As a result, some approaches [16, 17, 18] have begun exploring label-efficient strategies for SAM. WeSAM[16] and SlotSAM[18] leverage self-training techniques [19] with weak labels, such as points, boxes, and polygons, to generate pseudo-labels, allowing the network to predict complete masks. Cat-SAM [17] adopts a few-shot learning approach to fine-tune SAM using box prompts to predict masks. While using boxes or coarse masks as prompts have yielded excellent results, point annotations remain less effective than fully supervised methods. Moreover, point annotations are much cheaper than masks and boxes  [^1] https://cloud.google.com/ai-platform/data-labeling/pricing, particularly for RSIs with numerous dense objects. Therefore, this paper aims to fine-tune SAM for RSIs using the most challenging yet cost-effective point annotations.
@@
As shown in Fig. 1 (a), we first present the fine-tuning process of vanilla SAM using full mask annotations. It takes point or box prompts as input to generate the predicted mask $\mathcal{M}_{pred}$, which is supervised by the ground truth (GT) mask $\mathcal{M}_{gt}$. In contrast, the self-training-based method[16, 18] (depicted in Fig. 1 (b)) only requires pseudo-labels generated by the model itself. Specifically, the input undergoes both weak and strong augmentations separately and is fed into the network, resulting in $\mathcal{M}_{pred}$ and $\mathcal{M}^{\prime}_{pred}$, respectively. $\mathcal{M}_{pred}$ serves as a pseudo-label to constrain $\mathcal{M}^{\prime}_{pred}$, enabling iterative training. This method is feasible primarily due to the principles of source-free domain adaptation (SFDA)[20, 21, 22]. The core idea of SFDA is to improve model performance using unlabeled data from the target domain without requiring access to source domain data.
@@
However, self-training often depends on the quality of pseudo-labels. If there is noise in the pseudo-labels, the model may overfit incorrect patterns. To address this, two common approaches are feature alignment [23, 24, 25] and logit regularization [26, 16]. However, the former requires access to the distribution of source data, which is impractical for SAM. The latter can also affect results if the prediction of anchor logits is inaccurate. In contrast to these methods, our approach aligns the features of the source and target models at the image encoder. Rather than performing simple image-level alignment, we map the corresponding prompt locations to the encoder features for instance-level alignment. Since object point labels are already annotated, we do not rely on inaccurate predicted logits for constraints or use source data information. Specifically, before beginning self-training, we first extract features for each instance from the target data using the source model. We then cluster these instances using the parameter-free clustering algorithm FINCH [27] and compute target prototypes for all clusters. During self-training, we maintain a First-in-First-Out (FIFO) memory bank, which stores instance-level predicted features and computes the predicted prototypes similarly. Since discrepancies between the number of target and predicted prototypes may exist, direct correspondence cannot be established. To resolve this, we employ the Hungarian algorithm, which automatically matches these two types of prototypes and aligns them using a matching loss. We call this method Prototype-based Regularization (PBR).
@@
Refer to captionRefer to caption(a)(b)Negative promptPositive prompt
@@
Figure 2: (a) Segmentation results using only positive prompts. (b) Segmentation results using both positive and negative prompts.
@@
Moreover, RSIs are captured from overhead perspectives and contain densely detected objects and large-scale backgrounds, making points as prompts more semantically ambiguous because points lack boundary information. We tested RSIs on SAM’s demo website  [^2] https://segment-anything.com/demo; as shown in Fig. 2 (a), the densely distributed tennis courts in the image can cause the mask decoder to mistakenly interpret them as a single instance. However, after adding negative samples (shown in Fig. 2 (b)), the remaining parts were effectively removed. Inspired by this, selecting appropriate locations for negative prompts is crucial. To address this, we propose a method for adaptively extracting negative prompts during training, called Negative Prompt Calibration (NPC). The process is based on a prior assumption: there is no overlap between predicted masks of different instances. We first calculate the IoU between each instance and use other samples with an IoU above a certain threshold with respect to a given sample as candidate negative prompts. Then, we randomly select $k$ positive prompts to serve as negative prompts for the target sample. Finally, we input the new prompts into the mask decoder to obtain refined masks.
@@
We integrate the above two methods into the self-training-based point-supervised framework, named PointSAM. We conduct experiments on three representative RSI datasets: NWPU VHR-10, WHU, and HRSID. The results demonstrate that our approach effectively adapts vanilla SAM to various RSI scenarios under point supervision. Additionally, we apply PointSAM as a bounding box generator in point-supervised object detection tasks, indicating that this method can extend to other point-supervised applications. Our contributions are summarized as follows:
@@
  • We introduce Prototype-based Regularization (PBR), which aligns the features of source and target models at the instance level, utilizing dynamic prototype updating and the Hungarian algorithm to improve model generalization.
@@
  • We develop Negative Prompt Calibration (NPC), which adaptively adjusts negative prompts during training, enhancing the accuracy of predicted masks in dense scenarios.
@@
  • We demonstrate the effectiveness of PointSAM through extensive experiments on three RSI datasets (NWPU VHR-10, WHU, and HRSID), showing significant improvements in segmentation performance under point supervision. Additionally, we extend the application of PointSAM to bounding box generation in point-supervised oriented object detection tasks, showcasing its versatility and potential for broader use in point-based supervised learning scenarios.
@@ II-B Point-based supervision
Point annotations are often used to save on mask or box annotations. Compared to image-level annotations, it can indicate the object’s location, providing stronger priors for subsequent processing and offering better practicality. Point-supervised methods are widely applied in detection [38, 39, 40, 41, 42] or segmentation[43, 44, 16, 45, 46, 47] tasks. For example, P2BNet [39] uses Multiple Instance Learning (MIL) to select the box with the highest confidence from multiple boxes containing points. Point2Mask [44] formulates the pseudo-mask generation from points as an Optimal Transport (OT) problem. Unlike natural images, instances in RSIs are mostly smaller and more densely packed, making point annotations much more convenient for label generation. PointOBB [40] learns object scale and angle information through self-supervised learning across different views, enabling the generation of oriented bounding boxes from points. PMHO [41] first uses SAM as a point-to-mask converter. Then, it converts the initial mask into a horizontal bounding box (HBB) and uses an HBB-to-OBB network to obtain the final oriented bounding boxes (OBB). In our work, we aim to fine-tune the original SAM model using point annotations to better adapt it to RSIs. Consequently, a straightforward idea is to use the proposed PointSAM as a point-to-box converter, similar to PMHO. We also conducted experiments on weakly supervised oriented object detection and achieved promising results.
@@ II-D Recognition in Remote Sensing Images
Remote sensing images (RSIs) are captured by airborne or satellite sensors to observe and analyze the Earth’s surface. These images provide critical information for a wide range of applications, including environmental monitoring [53], urban planning [54], disaster management [55], and military operations [56]. A key characteristic of RSIs is their overhead perspective, typically categorized into optical images and Synthetic Aperture Radar (SAR) images. The objects detected in these images often exhibit significant scale variations and dense distributions. Existing methods address these challenges through feature processing [57, 58], loss function design [59, 60], and post-processing stages [61, 62]. However, these approaches are primarily tailored for object counting and detection tasks, with interactive segmentation remaining relatively unexplored. In our case, SAM often struggles to segment dense objects, especially when only points are used as prompts. If the positive prompt is not well-annotated, the predicted mask may become confused with the surrounding foreground. Negative prompts can help mitigate this issue, but selecting the correct negative prompt remains challenging. Therefore, we propose using network-adaptive learned negative prompts to calibrate the predicted masks.
@@ III-A2 Low-Rank Adaptation
Low-Rank Adaptation (LoRA) [36] is a technique used to reduce the computational and memory requirements of training large neural networks. By approximating weight updates with low-rank matrices, LoRA allows for more efficient fine-tuning of pre-trained models. This approach enables the adaptation of large models to new tasks or datasets with significantly lower resource consumption while maintaining performance. For each weight in the encoder network $\theta\in\mathbb{R}^{d_{i}\times d_{o}}$, we use a low-rank approximation $\omega=AB$ where $A\in\mathbb{R}^{d_{i}\times r}$ and $A\in\mathbb{R}^{r\times d_{o}}$ with $r$ indicating the rank. We can achieve a compression rate of $r\left(d_{i}+d_{o}\right)/d_{i}\cdot d_{o}$. Only $A$ and $B$ are updated via backpropagation during adaptation to reduce memory footprint. At the inference stage, the weight is reconstructed by combining the low-rank reconstruction and original weight, $\theta=\theta+AB$.
@@ III-B Pointly-supervised Segment Anything Model
The overall architecture of our proposed Pointly-supervised Segmentation Anything Model (PointSAM) is shown in Fig. 3. The pipeline is divided into two stages: Offline Prototype Generation and SAM with Self-Training. In the first stage, we extract prototypes of instances from the target dataset offline (see Sec. III-C1). During the self-training phase, two different views of each image, $\mathcal{I}_{img}$, are generated: $\mathcal{I}_{img}^{s}$ with strong data augmentation and $\mathcal{I}_{img}^{w}$ with weak data augmentation. The data augmentation techniques are detailed in [16]. Both $\mathcal{I}_{img}^{s}$ and $\mathcal{I}_{img}^{w}$ are then fed into shared teacher and student encoders. The encoder structures are fixed, and additional LoRA layers are used to fine-tune the model. The teacher’s image encoder outputs predicted instance features to a memory bank, which is updated using a First-in-First-Out (FIFO) strategy. These features are clustered to obtain new predicted prototypes (see Sec. III-C2). The offline-extracted prototypes are aligned with the new predicted prototypes using a matching loss (see Sec. III-C3). Simultaneously, the teacher and student networks predict corresponding masks $\mathcal{M}^{t}$ and $\mathcal{M}^{s}$. We apply the Negative Prompt Calibration (NPC) strategy to $\mathcal{M}^{t}$ to generate refined masks, $\mathcal{M}^{r}$ (see Sec. III-D), which are then used as pseudo-labels to train the student network. For more information on the network training loss, refer to Sec. III-E.

### v2

@@ Abstract
Segment Anything Model (SAM) is an advanced foundational model for image segmentation, which is gradually being applied to remote sensing images (RSIs). Due to the domain gap between RSIs and natural images, traditional methods typically use SAM as a source pre-trained model and fine-tune it with fully supervised masks. Unlike these methods, our work focuses on fine-tuning SAM using more convenient and challenging point annotations. Leveraging SAM’s zero-shot capability, we adopt a self-training framework that iteratively generates pseudo-labels. However, noisy labels in pseudo-labels can cause error accumulation. To address this, we introduce Prototype-based Regularization, where target prototypes are extracted from the dataset and matched to predicted prototypes using the Hungarian algorithm to guide learning in the correct direction. Additionally, RSIs have complex backgrounds and densely packed objects, making it possible for point prompts to mistakenly group multiple objects as one. To resolve this, we propose a Negative Prompt Calibration method, based on the non-overlapping nature of instance masks, where overlapping masks are used as negative signals to refine segmentation. Combining these techniques, we present a novel pointly-supervised segment anything model, PointSAM. We conduct experiments on three RSI datasets, including WHU, HRSID, and NWPU VHR-10, showing that our method significantly outperforms direct testing with SAM, SAM2, and other comparison methods. Additionally, PointSAM can act as a point-to-box converter for oriented object detection, achieving promising results and indicating its potential for other point-supervised tasks. The code is available at https://github.com/Lans1ng/PointSAM.
@@ I Introduction
Foundation models are versatile, large-scale models designed for a wide range of tasks and applications. They have demonstrated exceptional performance in areas such as natural language processing (e.g., BERT[1] and GPT-3[2]) and multimodal tasks (e.g., CLIP[3] and ALIGN [4]). Recently, Segment Anything Model (SAM) [5, 6] was introduced as a foundation model specifically for image segmentation. Trained on a billion-scale dataset of masks and prompts, SAM can be applied to various downstream tasks requiring promptable segmentation, including healthcare [7, 8], autonomous driving [9], and remote sensing [10, 11, 12].
Despite SAM’s strong zero-shot capabilities, challenges persist in handling out-of-distribution (OOD) data and domain shifts in remote sensing images (RSIs). Many categories in RSIs are not represented in SAM’s training data. Furthermore, RSIs are typically captured from aerial or satellite perspectives and differ significantly from natural images. Consequently, recent studies [10, 13, 14, 15] have focused on how to fine-tune SAM for specific tasks. For example, RS-Prompter [10] uses queries or anchors as prompts to guide SAM’s mask decoder for instance segmentation. Similarly, SAM-CD [13] employs FastSAM’s encoder and introduces adapters for fine-tuning in change detection tasks.
@@
While these methods achieve promising results, they rely on full mask annotations, which are difficult and time-consuming to obtain. To address this, recent approaches [16, 17, 18] have focused on label-efficient strategies for SAM. WeSAM[16] and SlotSAM[18] use self-training [19] with weak labels, such as points and boxes, to generate pseudo-labels, enabling the network to predict complete masks. Cat-SAM[17] adopts a few-shot learning approach, fine-tuning SAM with box prompts for mask prediction. While box and coarse mask prompts have shown strong results, point-based supervision remains less effective. Moreover, point annotations are much cheaper than masks and boxes  [^1] https://cloud.google.com/ai-platform/data-labeling/pricing, particularly for RSIs with numerous dense objects. Therefore, this paper aims to explore how to adapt SAM to RSIs using the most challenging yet cost-effective point annotations. First, let us review how full mask annotations are used to fine-tune SAM for RSIs. As shown in Fig. 1 (a), SAM takes point or box prompts as input to generate the predicted mask $\mathcal{M}_{pred}$, which is supervised by the ground truth (GT) mask $\mathcal{M}_{gt}$. In contrast, the self-training-based method[16, 18] (depicted in Fig. 1 (b)) only requires pseudo-labels generated by the model itself. Specifically, the input undergoes both weak and strong augmentations separately and is fed into the SAM’s image encoder, resulting in $\mathcal{M}_{pred}$ and $\mathcal{M}^{\prime}_{pred}$, respectively. $\mathcal{M}_{pred}$ serves as a pseudo-label to constrain $\mathcal{M}^{\prime}_{pred}$, enabling iterative training. This method is feasible primarily due to the principles of source-free domain adaptation (SFDA)[20, 21, 22]. The core idea of SFDA is to improve model performance using unlabeled data from the target domain without requiring access to source domain data.
@@
However, self-training often depends on the quality of pseudo-labels. If there is noise in the pseudo-labels, the model may overfit incorrect patterns. To address this, two common approaches are feature alignment [23, 24, 25] and logit regularization [26, 16]. However, the former requires access to the distribution of source data, which is impractical for SAM. The latter can also affect results if the prediction of anchor logits is inaccurate. In contrast to these methods, our approach aligns the features of the source and target models at the image encoder. Rather than performing simple image-level feature alignment, we map the corresponding prompt locations to the encoder features for instance-level feature alignment. Since object point labels are already annotated, we do not rely on inaccurate predicted logits for constraints or use source data information. Specifically, before beginning self-training, we first extract features for each instance from the target data using the source model. We then cluster these instances using the parameter-free clustering algorithm FINCH [27] and compute target prototypes for all clusters. During self-training, we maintain a First-in-First-Out (FIFO) memory bank, which stores instance-level predicted features and similarly computes the predicted prototypes. Since discrepancies between the number of targets and predicted prototypes may exist, direct correspondence cannot be established. To resolve this, we employ the Hungarian algorithm, which automatically matches these two types of prototypes and aligns them using a matching loss. We call this method Prototype-based Regularization (PBR).
@@
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption(a)(b)Negative promptPositive promptNWPU VHR-10WHUHRSIDPositive promptPositive promptNegative promptNegative prompt
@@
Figure 2: Segmentation results on the NWPU VHR-10, WHU, and HRSID datasets. (a) Segmentation results using only positive prompts. (b) Segmentation results using both positive and negative prompts.
@@
Moreover, RSIs are captured from overhead perspectives and contain densely detected objects and large-scale backgrounds, making points as prompts more semantically ambiguous because points lack boundary information. We tested RSIs on SAM’s demo website  [^2] https://segment-anything.com/demo; as shown in Fig. 2 (a), the densely distributed instances in the image can cause the mask decoder to mistakenly interpret them as a single instance. However, after adding negative samples (shown in Fig. 2 (b)), the remaining parts were effectively removed. Inspired by this, selecting appropriate locations for negative prompts is crucial. We thus propose a method for adaptively extracting negative prompts during training, called Negative Prompt Calibration (NPC). This procedure is based on a prior assumption: there is no overlap between predicted masks of different instances. We first calculate the IoU between each instance and use other samples with an IoU above a certain threshold with respect to a given sample as candidate negative prompts. Then, we randomly select $k$ positive prompts to serve as negative prompts for the target sample. Finally, we input the new prompts into the mask decoder to obtain refined masks.
@@
We integrate the above two methods into the self-training-based point-supervised framework, named PointSAM. We conduct experiments on three representative RSI datasets: NWPU VHR-10, WHU, and HRSID. The results demonstrate that our approach effectively adapts vanilla SAM to various RSI scenarios under point supervision. Additionally, we apply PointSAM as a bounding box generator in point-supervised oriented object detection tasks, indicating that this method can extend to other point-supervised applications. Our contributions are summarized as follows:
@@
  • We propose Prototype-based Regularization (PBR), which extracts instance-level features from both the source and target models. By using non-parametric clustering, dynamically updating prototypes, and Hungarian matching, PBR prevents the model from learning in incorrect directions and improves model generalization.
@@
  • We introduce Negative Prompt Calibration (NPC), which adaptively adjusts negative prompts during training. By using positive prompts from overlapping samples as negative prompts, NPC enhances the original SAM’s predicted masks and achieves more accurate results in dense scenarios.
@@
  • We are the first to leverage point annotations to fine-tune SAM for remote sensing images (RSIs). Extensive experiments on three datasets (NWPU VHR-10, WHU, and HRSID) validate the effectiveness of PointSAM, achieving significant improvements in point-supervised segmentation performance. Additionally, we extend PointSAM’s application to bounding box generation for point-supervised oriented object detection, demonstrating its versatility and potential in point-based supervised learning tasks.
@@ II-B Point-based supervision
Point annotations are often used to save on mask or box annotations. Compared to image-level annotations [38], it can indicate the object’s location, providing stronger priors for subsequent processing and offering better practicality. Point-supervised methods are widely applied in detection [39, 40, 41, 42, 43, 44] or segmentation[45, 46, 47, 48, 49, 50] tasks. For example, P2BNet [40] uses Multiple Instance Learning (MIL) to select the box with the highest confidence from multiple boxes containing points. Point2Mask [46] formulates the pseudo-mask generation from points as an Optimal Transport (OT) problem. Unlike natural images, instances in RSIs are mostly smaller and more densely packed, making point annotations much more convenient for label generation. PointOBB [41] learns object scale and angle information through self-supervised learning across different views, enabling the generation of oriented bounding boxes from points. PMHO [43] first uses SAM as a point-to-mask converter. Then, it converts the initial mask into a horizontal bounding box (HBB) and uses an HBB-to-OBB network to obtain the final oriented bounding boxes (OBB). In our work, we aim to fine-tune the original SAM model using point annotations to better adapt it to RSIs. Consequently, a straightforward idea is to use the proposed PointSAM as a point-to-box converter, similar to PMHO. We also conducted experiments on weakly supervised oriented object detection and achieved promising results.
@@ II-D Recognition in Remote Sensing Images
Remote sensing images (RSIs) are captured by airborne or satellite sensors to observe and analyze the Earth’s surface. These images provide critical information for a wide range of applications, including environmental monitoring [56], urban planning [57], disaster management [58], and military operations [59]. A key characteristic of RSIs is their overhead perspective, typically categorized into optical images and Synthetic Aperture Radar (SAR) images. The objects detected in these images often exhibit significant scale variations and dense distributions. Existing methods address these challenges through feature processing [60, 61, 62, 63], loss function design [64, 65, 66, 67], and post-processing stages [68, 69]. However, these approaches are primarily tailored for object counting and detection tasks, with interactive segmentation remaining relatively unexplored. In our case, SAM often struggles to segment dense objects, especially when only points are used as prompts. If the positive prompt is not well-annotated, the predicted mask may become confused with the surrounding foreground. Negative prompts can help mitigate this issue, but selecting the correct negative prompt remains challenging. Therefore, we propose using network-adaptive learned negative prompts to calibrate the predicted masks.
@@ III-A2 Low-Rank Adaptation
Low-Rank Adaptation (LoRA) [36] is a technique used to reduce the computational and memory requirements of training large neural networks. By approximating weight updates with low-rank matrices, LoRA allows for more efficient fine-tuning of pre-trained models. This approach enables the adaptation of large models to new tasks or datasets with significantly lower resource consumption while maintaining performance. For each weight in the encoder network $\theta\in\mathbb{R}^{d_{i}\times d_{o}}$, we use a low-rank approximation $\omega=AB$ where $A\in\mathbb{R}^{d_{i}\times r}$ and $B\in\mathbb{R}^{r\times d_{o}}$ with $r$ indicating the rank. We can achieve a compression rate $r=\left(d_{i}+d_{o}\right)/d_{i}\cdot d_{o}$. Only $A$ and $B$ are updated via backpropagation during adaptation to reduce memory footprint. At the inference stage, the weight is reconstructed by combining the low-rank reconstruction and original weight, $\theta=\theta+AB$.
@@ III-B Pointly-supervised Segment Anything Model
In our task, only point labels are available, and there is a significant domain gap between RSIs and natural images. Therefore, our approach focuses on addressing two key challenges: 1) how to effectively adapt SAM to RSIs, and 2) how to fully exploit the potential of point annotations. For the first challenge, we combine a self-training strategy with prototype representation constraints to ensure stable knowledge transfer and prevent the model from learning incorrect patterns. For the second challenge, we propose extracting features from point annotations to generate prototypes and refining mask predictions by adjusting negative point prompts.
The overall architecture of the Pointly-supervised Segment Anything Model (PointSAM) is illustrated in Fig. 3. The pipeline is divided into two main stages: Offline Prototype Generation and Self-Training with SAM. In the first stage, we extract instance prototypes offline from the target dataset (see Sec. III-C1). Target prototypes are generated through clustering and remain fixed during subsequent training. In the second stage, two augmented views of the input image $\mathcal{I}_{img}$ are generated: $\mathcal{I}_{img}^{s}$ with strong data augmentation and $\mathcal{I}_{img}^{w}$ with weak data augmentation [16]. These augmented images are then fed into shared teacher and student encoders. While the encoder structures remain frozen, the model is fine-tuned with additional LoRA layers. Since the teacher network is dynamically trained, its image encoder extracts instance features that are stored in a memory bank [71] updated using a First-in-First-Out (FIFO) strategy to ensure that stored features remain up-to-date. These stored features are clustered to generate new predicted prototypes (see Sec. III-C2). The target prototypes are aligned with the predicted prototypes using a Hungarian Matching loss (see Sec. III-C3). Meanwhile, the teacher and student networks generate corresponding masks, $\mathcal{M}^{t}$ and $\mathcal{M}^{s}$. For $\mathcal{M}^{t}$, we apply a Negative Prompt Calibration (NPC) strategy, using positive prompts from overlapping samples as negative prompts for specified samples to refine the mask predictions. This process results in optimized masks, $\mathcal{M}^{r}$ (see Sec. III-D), which are subsequently used as pseudo-labels to train the student network. For detailed information on the network training losses, refer to Sec. III-E.